OpenStudy (anonymous):

\[\log_{2} (8-6x)=5\]

OpenStudy (watchmath):

\(2^5=8-6x\) \(32-8=-6x\) \(24=-6x\) \(x=-4\)

OpenStudy (anonymous):

Thanks! :)